For y = f(x) = x - 2x + 4, find dy and Ay, given x = 5 and Ax = 0.2 dy = (Type an integer or a decimal.)
Answer #1

- - d da عهعسهرکهولا122 = (د) عن 4 = 002 < ۵ بعدی. په ده(2- ا = وه وقد علم ،) = (22) عق= ه by z dy 2.م = لا -1) = 0.2 (7 کد2.s ਜs - ਹਾਦਸਾ tigor he - - ( Cਹਤਮੰਦ